OpenStudy (anonymous):

are the following functions equal? y+1=7(x+2) y=7x+13 A. yes the functions are equal because they have the same domain though they do not have the same function rule. B. yes the functions are equal because they have the same function even though they do not have the same domain C. no, the functions are not equal because they don't have the same function rule D. yes, the functions are equal because they have the same domains and same function rule. E. no, the functions are not equal because they do not have the same domain.

OpenStudy (anonymous):

D \[y+1=7(x+2) \]multiply x+2 by 7 \[y+1=7x+14\]subtract the 1 and it becomes \[y=7x+13\]

OpenStudy (anonymous):

okay thank you!